OpenStudy (anonymous):

logx-log(x+4)=-1 log(x/(x+4))=-1 ?

jimthompson5910 (jim_thompson5910):

So far, so good Now convert that logarithmic equation to exponential form You do this by using the rule log(x) = y ----> x = 10^y --------------------------------------------------------- So that means log(x/(x+4))=-1 turns into x/(x+4) = 10^(-1) x/(x+4) = 1/(10^1) x/(x+4) = 1/10 and I'll let you finish up

OpenStudy (anonymous):

x=x+4/10 ? That's actually where I was, I should have gone further, because it's that point where I don't know how to get the variable on one side

jimthompson5910 (jim_thompson5910):

From here, you cross multiply x/(x+4) = 1/10 10x = 1(x+4) 10x = x + 4 10x - x = 4 9x = 4 x = 4/9

jimthompson5910 (jim_thompson5910):

Don't forget to check all possible answers. So plug x = 4/9 back into the original equation and simplify. You should get an identity.

OpenStudy (anonymous):

ahhhhh.